OpenStudy (anonymous):

The Scott family and the Phillips family each used their sprinklers last summer. The water output rate for the Scott family's sprinkler was 30L per hour. The water output rate for the Phillips family's sprinkler was 20L per hour. The families used their sprinklers for a combined total of 50 hours, resulting in a total water output of1300L. How long was each sprinkler used? Scott family: hours Phillips family: hours Will medal

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

write and solve a system of equations

OpenStudy (anonymous):

I dont know how to do that

Vocaloid (vocaloid):

start by defining two variables let s = number of hours used for the Scott family let p = number of hours used by the Phillips family

OpenStudy (anonymous):

yea

Vocaloid (vocaloid):

p + s = 50

OpenStudy (anonymous):

ok

Vocaloid (vocaloid):

20p + 30s = 1300 p + s = 50 solve for p and s

OpenStudy (anonymous):

thats the part im stuck on

Vocaloid (vocaloid):

start by multiplying the second equation by 20

OpenStudy (anonymous):

which one

Vocaloid (vocaloid):

I said the second one

OpenStudy (anonymous):

doesnt make any sense

Vocaloid (vocaloid):

here are your two equations 20p + 30s = 1300 p + s = 50 multiply the second equation by 20

OpenStudy (anonymous):

20p+20s=100

Vocaloid (vocaloid):

almost

Vocaloid (vocaloid):

50*20 = ?

OpenStudy (anonymous):

1000

Vocaloid (vocaloid):

good.

Vocaloid (vocaloid):

now we have 20p + 30s = 1300 20p + 20s = 1000 now subtract the two equations.

OpenStudy (anonymous):

10s=300

Vocaloid (vocaloid):

good now find s...

OpenStudy (anonymous):

s=30

Vocaloid (vocaloid):

great, now go back and find p

Vocaloid (vocaloid):

we know that p + s = 50, so what is p?

OpenStudy (anonymous):

25

Vocaloid (vocaloid):

nope.

Vocaloid (vocaloid):

s = 30 p = ?

OpenStudy (anonymous):

20

Vocaloid (vocaloid):

great.

OpenStudy (anonymous):

ty have a few more up to helping me